# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

### v6

@@ Abstract
The application of formulas (e.g., physics formulas) is a fundamental human ability in solving numerical reasoning problems. Existing numerical reasoning datasets rarely explicitly state the formulas employed, as their questions often rely on implicit commonsense mathematical knowledge. To address this gap, we introduce FormulaReasoning, a new dataset specifically designed for formula-based numerical reasoning. It consists of 5,324 questions that require numerical calculations grounded in external physics formulas. We provide normalized, fine-grained annotations in both English and Chinese, including formula structures, parameter names, symbols, numerical values, and units—curated through extensive manual effort with LLM-assisted validation to ensure high quality. Additionally, we offer a consolidated formula database to serve as an external knowledge source. We analyze various reasoning approaches on FormulaReasoning, with emphasis on comparative evaluation of different architectural and methodological frameworks. Our assessment includes retrieval-augmented methods, approaches that decompose reasoning into formula generation, parameter extraction, and numerical calculation, as well as optimization techniques using preference data. We identify key challenges in formula-based numerical reasoning that require further investigation across different reasoning paradigms, highlighting opportunities for methodological advancement.
@@ 1 Introduction
Take a question from the GSM8K dataset (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?” This question implicitly uses simple commonsense mathematical knowledge (e.g., half means dividing by 2) to solve without requiring complex domain-specific formular knowledge for numerical reasoning. Recently, Liu et al. Liu et al. (2023) constructed two formula-based datasets, Math23K-F and MAWPS-F. Although 33% and 38% of the questions in these datasets, respectively, require the use of formulas, they still mainly consist of simple commonsense formulas (e.g., total_amount = unit_amount $\times$ total_number). Currently, there is a lack of datasets where questions require domain-specific formulas to guide complex numerical reasoning, such as the physics formula used to calculate the heat absorption of an object.
@@
To fill the gap, we constructed FormulaReasoning, a new dataset for numerical reasoning that requires the use of domain-specific formulas (i.e., physics). We annotated formulas for each question as illustrated in Figure 1. This formula-based feature enables FormulaReasoning to effectively differentiate model capabilities across different scales, particularly highlighting the limitations of smaller language models in handling domain-specific formula applications. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known formula-free datasets GSM8K and MATH (Hendrycks et al., 2021). Compared to Math23K-F and MAWPS-F, FormulaReasoning contains more diverse formulas (272 vs. 18–51). Moreover, the higher average number of reasoning steps (2.39 vs. 1.01–1.16) provides a more comprehensive evaluation framework for understanding how different model architectures handle multi-step formula-based reasoning.
Table 1: Statistics of Math23-F, MAWPS-F, GSM8K, MATH, and our FormulaReasoning.
@@
We conducted extensive experiments on FormulaReasoning to compare different reasoning paradigms, including standard LLM approaches with 8B to >>100B parameters, fine-tuned models with Chain-of-Thought (CoT) supervision, data augmentation methods, and retrieval-augmented generation (RAG) methods. In addition, we derived preference data and evaluated direct preference optimization (DPO) methods. Our analysis reveals significant differences in how various approaches handle formula-based reasoning.
@@
FormulaReasoning is available from HuggingFace https://huggingface.co/datasets/cat-overflow/FormulaReasoning. Our code is available from GitHub https://github.com/nju-websoft/FormulaReasoning.
2 Related Work
2.1 Numerical Reasoning Datasets
2.2 Numerical Reasoning Methods
Our experiments implemented and evaluated representative existing methods as baselines for FormulaReasoning, including zero-shot CoT prompting for LLMs ranging from 8B to more than 100B parameters. We trained a specialized formula retriever to be used in RAG. We divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned small models. We derived preference data and performed DPO.
@@ 3.2 Formula Normalization
We checked the correctness of the formula format in the explanations by rules, including whether there were omissions in parameter names, symbols, or units, and these minor issues were correctable. To assess the correctness of each normalized explanation, we extracted formulas from the explanation and calculated an answer using the Numbat calculator [^2] https://numbat.dev. Numbat is designed for scientific computation with support for physical units.. If we detected that the LLM generated an ill-formatted or incorrect normalized explanation, we used a few-shot prompt to identify and correct these specific errors. We observed that the questions that remained incorrect despite multiple attempts were of inherently poor quality, e.g., missing important reasoning steps, as illustrated in Appendix A.1.2. These questions were removed from our dataset. After that, our dataset contains 5,420 questions remaining.
@@ 3.5 English Version of FormulaReasoning
We employed the LLM to translate Chinese questions, explanations, and formula database into English. The translation prompts are provided in Appendix A.1.4. To evaluate translation quality, we manually translated a sample of 50 questions and calculated the perplexity of the LLM-translated version and the manually translated version using Qwen models for comparison. We computed the perplexity of LLM and human translations using the Qwen2.5-14B-Instruct model, resulting in scores of $6.28$ and $7.26$, respectively. These findings demonstrate that the current English version maintains a reasonable level of fluency.
We also manually compared the Chinese and translated English versions of the 50 sampled questions. We found that no question had translation errors that would affect question answering.
We provide both Chinese and English versions of FormulaReasoning.
@@ 4 Experimental Setup
We implemented a wide range of reasoning approaches for handling the questions within FormulaReasoning, with a focus on understanding how different methodological choices impact performance on formula-based numerical reasoning tasks. Our investigation includes analyzing various prompting strategies for language models, developing specialized fine-tuning techniques, and incorporating external knowledge through retrieval-augmented methods. Specifically, we developed two complementary approaches to enhance reasoning capabilities: one that decomposed the reasoning process into formula generation, parameter extraction, and numerical calculation, and another that leveraged data augmentation techniques. We also explored preference learning to refine reasoning processes. This comprehensive experimental setup allows us to compare the strengths and limitations of different architectural and methodological choices for formula-based reasoning.
@@ 4.1 Dataset Split
We divided FormulaReasoning into three subsets: training, HoF (Homologous Formulas) test, and HeF (Heterologous Formulas) test, comprising 4,524, 413, and 387 questions, respectively. All formulas in the HoF test set appeared in the training set, while in the HeF test set, each question required at least one formula not seen in the training set. This division was to evaluate the generalizability of fine-tuned models on new formulas.
@@ 4.2 Evaluated Methods
We conducted experiments across a diverse spectrum of models including cost friendly small-scale models including GLM-4.5-Air (GLM et al., 2024), ERNIE-4.5-21B-A3B Baidu-ERNIE-Team (2025), Gemma-3N-E4B Gemma (2025), Llama-3.3-8B-Instruct (Meta, 2024), Mistral-Small-3.2-24b-Instruct (Mistral, 2025), GPT-oss-20B (Agarwal et al., 2025), Gemini-2.5-Flash-Lite (Comanici et al., 2025) and Seed-OSS-36B-Instruct Team (2025), alongside high-cost LLMs such as GPT-4o (OpenAI, 2024), GLM-4-plus (GLM et al., 2024), Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025), o3-mini (OpenAI, 2025b), and GPT-5 (OpenAI, 2025a). This comprehensive evaluation allowed us to assess capabilities across different model scales while identifying persistent challenges in formula reasoning.
@@
We found that directly prompting models having fewer than 8B parameters did not produce satisfactory results, so we performed supervised fine-tuning of small models. We conducted experiments on Qwen2.5-Math-7B-Instruct and Qwen2.5-Math-1.5B-Instruct.
Unlike large models, small models struggled with numerical extraction and calculation. To enhance their reasoning capabilities, we developed:
  • CoT-Supervised Fine-Tuning (CoT-SFT): A three-step process where the model generates relevant formulas, extracts parameter values and units from questions, and uses Numbat for final calculation.
  • Data Augmentation (DA): Using Qwen-max to generate and verify new training examples with correct reasoning steps. Details are provided in Appendix A.2.3.
@@ 4.3 Evaluation Metrics
We utilized Numbat to evaluate the prediction generated by each model against the gold-standard answer. A prediction was deemed correct if the relative error, (prediction - gold) / gold, was less than 1%. We used accuracy, which is the proportion of questions correctly answered, as our metric.
We also used PRM score to evaluate the reasoning steps generated by LLMs. There were two evaluation settings: one-step and multi-step. One-step evaluation scored the overall output using PRM, whereas multi-step evaluation divided the output into multiple steps, scored each individually, and took the average as the final score. Following the setting of DPO mentioned above, here we again used Qwen2.5-Math-PRM-7B as the PRM.
@@ 5 Experimental Results
We present the evaluation results on the Chinese version of FormulaReasoning. The results on the English version are in Appendix A.3.1, which are consistent with those on the Chinese version. The results of PoT are in A.3.2, not better than CoT.
@@
Model HoF test HeF test Avg. GLM-4.5-Air 59.32 61.50 60.37 ERNIE-4.5-21B-A3B 66.34 67.96 67.12 Gemma-3n-E4B-it 14.04 14.47 14.25 Llama-3.3-8b-Instruct 8.47 9.56 9.00 Mistral-Small-3.2-24b-Instruct 74.33 72.35 73.37 GPT-oss-20B 85.96 82.95 84.50 Gemini-2.5-Flash-Lite 81.11 81.40 81.25 Seed-OSS-36B-Instruct 81.84 84.24 83.00 GPT-4o 69.49 65.37 67.50 GLM-4-plus 61.74 63.57 62.63 Qwen3-235B-A22B-Thinking-2507 80.63 83.98 82.25 DeepSeek-R1 88.62 91.47 90.00 o3-mini 92.01 89.66 90.87 GPT-5 92.01 92.76 92.37 Human 93.49 90.47 92.03
@@ 5.1 Results of LLMs with CoT Prompts
We evaluated a diverse range of LLMs across different parameter scales on FormulaReasoning. Our analysis reveals that model performance correlates with both parameter scale and architectural design. As shown in Table 4, large-scale models (e.g., DeepSeek-R1 and o3-mini) achieved competitive results with accuracies exceeding 88% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Llama-3.3-8B and Gemma-3N-E4B-it exhibit accuracies below 15%, while even mid-sized models like GLM-4-Air only reached around 60%. Notably, despite their impressive capabilities, even advanced large-scale models such as GPT-4o (67.50%), GLM-4-plus (62.63%), and Qwen3-235B (82.25%) still trailed behind human performance (92.03%). This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
@@ 5.2 Results of LLMs with Formula Retriever
We found that incorporating retrieved formulas into the prompts improved the performance of LLMs. For GLM-4-Air, the accuracy increased by 1.94 from 65.13 to 67.07 on the HoF. Note that it would be meaningless to carry out this experiment on the HeF test set where the required formulas were not included when training the retriever.
@@ 5.3 Results of Supervised Fine-Tuned Models
The results of supervised fine-tuned small models are included in Appendix A.3.4. Both CoT-SFT and DA helped improve accuracy. In particular, Qwen2.5-Math-7B-Instruct with CoT-SFT outperformed GPT-4o, demonstrating the effectiveness of our fine-tuning method: delegating numerical calculations to a calculator and focusing on CoT reasoning.

### v7

@@ Abstract
The application of physics formulas is a fundamental human capability in numerical reasoning. While existing datasets often rely on implicit mathematical knowledge, they rarely explicitate the underlying formulas. To address this, we introduce FormulaReasoning, a new benchmark for formula-based numerical reasoning comprising 5,324 questions requiring calculations grounded in external physics principles. We provide high-quality, fine-grained annotations in English and Chinese—including formula structures, parameter names, symbols, values, and units—curated through manual effort and LLM-assisted validation. Additionally, we provide a consolidated formula database as an external knowledge source. To further challenge model performance, we develop an extended version of the dataset by coupling multiple questions. We evaluate various architectural and methodological frameworks, including retrieval-augmented methods, modular reasoning (formula generation, parameter extraction, and calculation), and preference-based optimization. Our analysis identifies critical challenges in formula-based reasoning, highlighting significant opportunities for future methodological advancement.
@@ 1 Introduction
Table 1: Statistics of Math23K-F, MAWPS-F, GSM8K, MATH, and our FormulaReasoning.
@@
Take a question from the GSM8K dataset (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?” This question implicitly uses simple commonsense mathematical knowledge (e.g., half means dividing by 2) to solve without requiring complex domain-specific formular knowledge for numerical reasoning. Recently, Liu et al. Liu et al. (2023) constructed two formula-based datasets, Math23K-F and MAWPS-F. Although 33% and 38% of the questions in these datasets, respectively, require the use of formulas, they still mainly consist of simple commonsense formulas (e.g., total_amount = unit_amount $\times$ total_number). Currently, there is a lack of datasets where questions require domain-specific formulas to guide complex numerical reasoning, such as the physics formula used to calculate the heat absorption of an object. Related work is in Appendix A.1.
@@
To fill this gap, we introduce FormulaReasoning, a dataset for numerical reasoning requiring domain-specific (physics) formulas. We provide fine-grained annotations (Figure 1), enabling effective differentiation of model capabilities, especially for smaller models. Table 1 compares FormulaReasoning with existing datasets. Compared to Math23K-F and MAWPS-F, FormulaReasoning contains more diverse formulas (272 vs. 18–51). Moreover, the higher average number of reasoning steps (2.36 vs. 1.01–1.16) provides a more comprehensive evaluation framework for understanding how different model architectures handle multi-step formula-based reasoning.
@@
We evaluated FormulaReasoning using various reasoning paradigms: standard LLMs (4B to >>100B parameters), fine-tuned models with CoT supervision, data augmentation, and RAG. We also derived preference data for DPO. Our analysis reveals significant differences in how approaches handle formula-based reasoning.
To further challenge the reasoning capabilities of models, we constructed FormulaReasoning+. This extended version coupled each question with another question and reformulated direct calculation as solving a system of equations, significantly increasing the difficulty of numerical reasoning.
Broader impact
It is worth noting that while FormulaReasoning is constructed within the physics domain, the underlying framework is domain-agnostic. By abstracting formulas into an operator-parameter structure, our dataset construction pipeline can be readily extended to other disciplines requiring structured knowledge, such as finance (e.g., compound interest formulas), chemistry (e.g., stoichiometry), and geometry. Furthermore, beyond serving as a benchmark, FormulaReasoning offers significant research value for step-wise reasoning, structured formula retrieval, and preference optimization (DPO), providing rich resources for advancing neuro-symbolic reasoning.
@@
FormulaReasoning is available from anonymous GitHub https://github.com/nju-websoft/FormulaReasoning under Apache 2.0.
@@ 2.2 Formula Normalization
We checked the correctness of the formula format in the explanations by rules, including whether there were omissions in parameter names, symbols, or units, and these minor issues were correctable. To assess the correctness of each normalized explanation, we extracted formulas from the explanation and calculated an answer using the Numbat calculator [^2] https://numbat.dev. Numbat is designed for scientific computation with support for physical units.. We achieved a formula normalization accuracy of 85.95% through this programmatic verification. We used few-shot prompts to correct LLM errors and removed poor-quality questions, such as those missing reasoning steps, as illustrated in Appendix A.2.2. After that, our dataset contains 5,420 questions remaining. To further improve data quality, we performed a rigorous human audit on a critical subset of the data. Specifically, we manually reviewed and corrected 363 questions (approximately 45.4% of the test set; see Section 3.1 for dataset split), which primarily consisted of questions where the model predictions deviated from the ground truth.
@@ 2.5 English Version of FormulaReasoning
LLMs translated the Chinese questions, explanations, and formulas into English using prompts from Appendix A.2.4. All components, including parameter names and reasoning steps, were translated.
To assess translation quality, we compared the perplexity of LLM and manual translations for 50 sampled questions using the Qwen2.5-14B-Instruct model. The LLM scored $6.28$ compared to $7.26$ for human versions, demonstrating reasonable fluency. Additionally, a manual review of these 50 questions found no translation errors that would influence consistency or question-answering accuracy.
2.6 Question Coupling
To further elevate the reasoning complexity, we introduced an extended version FormulaReasoning+ that tightly coupled the original questions with a synthesized auxiliary question and reformulated direct calculation as solving a system of equations. First, we selected a mutable parameter in the original question to mask as $[X]$ and denoted the original answer as $[Y]$. Second, we constructed a linear expression $E(X,Y)$ (e.g., $c_{1}Y+c_{2}X$) ensuring a positive result and magnitude alignment between $X$ and $Y$ by adjusting coefficients. Third, we calculated the value $v=E(X,Y)$ and prompted an LLM to generate an independent physics question ($Q_{aux}$) whose answer is $v$. Finally, we formulated a hybrid question that requires first solving $Q_{aux}$ to obtain $v$ and then solving a system of equation in $X$ and $Y$ where one equation represents the original question and the other is $v=E(X,Y)$. After removing the questions with failed coupling, the final FormulaReasoning+ contains 4,392 extended questions. It is worth noting that Question Coupling and Question Alteration (Section 2.4) are parallel processes, both derived from the same original questions. For the detailed coupling process, please refer to Appendix A.2.5.
@@ 3 Experimental Setup
To study how methodological choices affect formula-based reasoning in FormulaReasoning, we analyzed prompting, fine-tuning, and retrieval-augmented methods. We developed two specific approaches: one decomposing reasoning into formula generation and calculation steps, and another utilizing data augmentation. We also explored preference learning for refinement. This setup enables a comprehensive comparison of architectural and methodological impacts on performance.
@@ 3.1 Dataset Split
We divided FormulaReasoning into three subsets: training, HoF (Homologous Formulas) test, and HeF (Heterologous Formulas) test, comprising 4,524, 413, and 387 questions, respectively. Inheriting the same split from FormulaReasoning, FormulaReasoning+ contains 3,608 training questions, with the HoF test and HeF test sets containing 406 and 378 questions, respectively. All formulas in the HoF test set appeared in the training set, while in the HeF test set, each question required at least one formula not seen in the training set. This division was to evaluate the generalizability of fine-tuned models on new formulas.
@@ 3.2 Evaluated Methods
Human Performance
We recruited 108 students from a high school, each student being assigned 7–8 questions. Each student was given 40 minutes to complete these questions. These questions were used as part of their in-class exercises and, at the end, each student received a gift. The final statistics were collected to assess human performance, which was consented to by all students.
@@
We conducted experiments across a diverse spectrum of models including both cost-friendly small-scale models and high-cost LLMs. The list of evaluated models is provided in Appendix A.3.1. This comprehensive evaluation allowed us to assess capabilities across different model scales while identifying persistent challenges in formula reasoning.
@@
We found that directly prompting models having fewer than 8B parameters did not produce satisfactory results, so we performed supervised fine-tuning of small models (Qwen2.5-Math-7B-Instruct and Qwen2.5-Math-1.5B-Instruct). Unlike large models, small models struggled with numerical extraction and calculation. To enhance their reasoning capabilities, we developed: (1) CoT-Supervised Fine-Tuning (CoT-SFT), a three-step process where the model generates relevant formulas, extracts parameter values and units from questions, and uses Numbat for final calculation; and (2) Data Augmentation (DA), using Qwen-max to generate and verify new training examples with correct reasoning steps (details in Appendix A.3.4).
@@ 3.3 Evaluation Metrics
We utilized Numbat to evaluate the prediction generated by each model against the gold-standard answer. A prediction was deemed correct if the relative error, (prediction - gold) / gold, was less than 1%. We used accuracy, which is the proportion of questions correctly answered, as our metric. Crucially, our evaluation leveraged Numbat for dimensional analysis, allowing us to automate the unit conversion (e.g., $1000\text{ J}=1\times 10^{3}\text{ N}\cdot\text{m}$) beyond simple numerical matching.
To evaluate the quality of the complete reasoning steps, we also used PRM score to evaluate the reasoning steps generated by LLMs. There were two evaluation settings: one-step and multi-step. One-step evaluation scored the overall output using PRM, whereas multi-step evaluation divided the output into multiple steps, scored each individually, and took the average as the final score. Following the setting of DPO mentioned above, here we again used Qwen2.5-Math-PRM-7B as the PRM. We adopted PRM scoring instead of strict formula matching because valid reasoning paths in formula-based problems are often non-unique; different formula combinations can lead to the correct answer.
@@ 4 Experimental Results
We present the evaluation results on the Chinese version of FormulaReasoning. The results on FormulaReasoning+ are in Table 5. The results on the English version are in Appendix A.4.1, which are consistent with those on the Chinese version. The results of PoT are in A.4.2, not better than CoT.
@@
Model HoF HeF Avg. ERNIE-4.5-21B-A3B 66.34 67.96 67.12 Gemma-3n-E4B-it 14.04 14.47 14.25 GPT-oss-20B 85.96 82.95 84.50 GPT-4o 69.49 65.37 67.50 GLM-4-plus 61.74 63.57 62.63 Qwen3-235B-A22B 80.63 83.98 82.25 DeepSeek-R1 88.62 91.47 90.00 o3-mini 92.01 89.66 90.87 GPT-5 92.01 92.76 92.37 Human 93.49 90.47 92.03
Table 5: Accuracy of LLMs with CoT prompts on FormulaReasoning+.
Model HoF HeF Avg. ERNIE-4.5-21B-A3B 6.89 6.08 6.48 Gemma-3n-E4B-it 0.99 1.32 1.16 GPT-oss-20B 59.36 57.41 58.38 GPT-4o 15.27 17.72 16.50 GLM-4-plus 17.98 16.67 17.32 Qwen3-235B-A22B 78.08 77.78 77.93 DeepSeek-R1 83.99 80.42 82.20 o3-mini 73.89 73.96 73.93 GPT-5 76.11 73.28 74.69
@@ 4.1 Results of LLMs with CoT Prompts
We evaluated a diverse range of LLMs, as shown in Table 4 and Table 5. Our analysis reveals that model performance correlates with both parameter scale and architectural design. As shown in Table 4, large-scale models (e.g., DeepSeek-R1 and o3-mini) achieved competitive results with accuracies exceeding 88% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Gemma-3N-E4B-it exhibit accuracies below 15%. Notably, despite their impressive capabilities, even advanced large-scale models such as GPT-4o (67.50%), GLM-4-plus (62.63%), and Qwen3-235B-A22B (82.25%) still trailed behind human performance (92.03%). This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
The results on FormulaReasoning+, as shown in Table 5, demonstrate the increased difficulty of the extended dataset. All models experienced a drop in performance. Notably, the accuracy of GPT-4o and GLM-4-plus declined sharply to 16.50% and 17.32% respectively. Even GPT-5 showed a decrease of 17.68%, highlighting the challenges posed by the more complex reasoning tasks.
@@ 4.2 Results of LLMs with Formula Retriever
We found that incorporating retrieved formulas into the prompts improved the performance of LLMs. For example, for ERNIE-4.5-21B-A3B, the accuracy increased by 0.72 from 66.34 to 67.06 on the HoF. Note that it would be meaningless to carry out this experiment on the HeF test set where the required formulas were not included when training the retriever.
@@
Table 6: Accuracy of supervised fine-tuned models.
Model HoF HeF Avg. Qwen2.5-Math-7B-Instruct 65.62 58.91 62.27    + CoT-SFT 72.64 +7.02+7.02 65.63 +6.72+6.72 69.14 +6.87+6.87    + DA 67.55 +1.93+1.93 62.53 +3.62+3.62 65.04 +2.77+2.77 Qwen2.5-Math-1.5B-Instruct 64.41 55.56 59.99    + CoT-SFT 66.34 +1.93+1.93 53.75 1.81-1.81 60.05 +0.06+0.06    + DA 66.59 +2.18+2.18 59.17 +3.61+3.61 62.88 +2.89+2.89
@@ 4.3 Results of Supervised Fine-Tuned Models
In particular, Qwen2.5-Math-7B-Instruct with CoT-SFT outperformed GPT-4o on FormulaReasoning, demonstrating the effectiveness of our fine-tuning method: delegating numerical calculations to a calculator and focusing on CoT reasoning, although such a comparison might not be entirely fair. Data augmentation also improved the reasoning capacity of small models.